OpenStudy (anonymous):

Solve the following equation, giving the exact solutions which lie in [0, 2π). (Enter your answers as a comma-separated list.) tan^2(x)=3/2sec(x)

OpenStudy (anonymous):

can u plz rewrite ur question more explicitly...

OpenStudy (anonymous):

\[\tan ^{2}(x)=\frac{ 3 }{ 2 }\sec(x)\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

Change them into \(\sin\) and \(\cos\).

OpenStudy (anonymous):

now its bit more clear

OpenStudy (anonymous):

thus we have sin^2(x)/cos^2(x) =(3/2cosx) or (1 - cos^2(x))/cos^2(x) =3/(2cosx) ... i think now it can be solved after cross multiplication)

OpenStudy (anonymous):

matri could you go on? Im kinda clueless

OpenStudy (anonymous):

where do you find the prob in the above expression

OpenStudy (anonymous):

how do i cross multiply that?

OpenStudy (anonymous):

as u usually do it for regular equation

OpenStudy (anonymous):

2cosx *((1 - cos^2(x)) = 3 cos^2(x) .....

OpenStudy (anonymous):

2cosx *(1 - cos^2(x)) - 3 cos^2(x) =0 or cosx*(2(1 - cos^2(x)) -3cosx)=0...... cosx*(2(1 - cos^2(x) )-3cosx) =0 cosx*(2 - 2cos^2(x) -3cosx)=0 cosx*(2 -4cosx +cosx -2cos^2(x) )=0 cosx * (2*(1-2cosx) +cosx(1-2cosx))=0 or cosx *(2+cosx)*(1-2cosx)=0 hence cosx = 0 or cosx =1/2 [as cosx <>-2]

OpenStudy (anonymous):

i hope further u can do it...